OpenStudy (bacon):

Please Help!! How do i find y in... 18*(4y/3)+15y=1026?

OpenStudy (coolshubhs):

18*4y/3 +15y=1026 72y/3 + 15y =1026 24y+15y=1026 31y=1026 y=1026/31